# VLMs have Tunnel Vision: Evaluating Nonlocal Visual Reasoning in Leading VLMs

### v1

@@ Abstract
Visual Language Models (VLMs) excel at complex visual tasks such as VQA and chart understanding, yet recent work suggests they struggle with simple perceptual tests. We present an evaluation that tests vision-language models’ capacity for nonlocal visual reasoning- reasoning that requires chaining evidence collected from multiple, possibly distant, regions of an image. We isolate three distinct forms of non-local vision: comparative perception, which demands holding two images in working memory and comparing them; saccadic search, which requires making discrete, evidence-driven jumps to locate successive targets; and smooth visual search, which involves searching smoothly along a continuous contour. Flagship models (e.g., Gemini 2.5 Pro, Claude Vision 3.7, GPT-o4-mini), even those that perform well on prior primitive-vision benchmarks, fail these tests and barely exceed random accuracy on two variants of our tasks that are trivial for humans. Our structured evaluation suite allows us if VLMs can perform similar visual algorithms to humans. Our findings show that despite gains in raw visual acuity, current models lack core visual reasoning capabilities.
@@ 1 Introduction
Vision-Language Models (VLMs) have demonstrated impressive performance on complex multimodal tasks. These models appear to possess a deep understanding of images and achieve over 90% accuracy on benchmarks like AI2D and ChartQA[13, 15]. Such high-level competence would suggest these models have strong primitive visual skills. However, recent work such as VLMs are Blind[19] or HallusionBench [7], shows that VLMs excel at answering high-dimensional questions that require background context but fail at recognizing simple geometry. Newer models score higher on these adversarial benchmarks, which indicates they possess stronger visual acuity.
@@
This strategy works only when combined with known assumptions about an image. Graph comprehension questions are embedded with strong priors, such as standardized layouts and predictable elements (e.g., axes, legends), and may not require the model to reason over nonlocal regions. Models could instead learn to exploit these learned commonalities and only superficially parse graphs. This reliance on convention over direct visual evidence would stunt models from parsing images that defy these norms. Indeed, recent literature supports the hypothesis that VLMs are not as proficient at reading graphs as benchmark scores suggest [16]. For example, Masry et al. [16] showed that top models’ performance catastrophically declined on their novel ChartQAPro dataset compared to ChartQA, indicating a missing skill set for handling diverse, unseen charts. This, along with findings from benchmarks like HallusionBench [7], suggests VLMs still prioritize background knowledge over the visual evidence presented to them.
To address these deficiencies and probe the underlying algorithmic visual skills necessary for robust image understanding, we identify three core types of nonlocal visual reasoning. First, comparative perception is the qualitative comparison of visual entities even when precise discrepancies are difficult to articulate (e.g., recognizing that two complex shapes are not identical without explicitly itemizing every differing feature). Second, saccadic search, named after the rapid eye movements humans make, is the process of gathering and integrating information from different image regions. Each piece of evidence informs the next step; for example, this process is used when consulting a chart’s legend, then locating the corresponding line, then referencing an axis. Third, smooth visual search describes the continuous tracing of visual elements, such as following the outline of an object or tracing a curve to its conclusion. To systematically evaluate these capabilities, we introduce a procedurally-generated evaluation set comprising three task categories designed to be trivial for humans and require minimal prior knowledge. We present three task categories: Object Re-Identification, Visual Scavenger Hunt, and Circuit Connections. Examples of these tasks appear in Figure 1.
@@
    When are state-of-the-art Vision Language Models (VLMs) susceptible to errors in basic perception? In nonlocal visual reasoning tasks, do these initial perceptual errors compound or self-correct?
@@
    Can VLMs perform comparative perception and saccadic search? If so, must these models use natural language judgments to guide these processes, or can they execute these tasks through more direct visual analysis?
@@
All tested models perform poorly on at least one variant of every task that is trivial for humans. The performance gradient on Object Re-Identification shows that modern models selectively choose when to examine an image closely. Our work further shows that fuzzy vision interferes with visual reasoning and the models seem unable to self-correct. They instead rely on their prior natural language judgments over direct evidence in the image. VLMs struggle the most with smooth visual search and attempt to use other visual cues in the image to compensate for their lack of ability.
@@
  • We present three procedural generators for an evaluation suite. Each generator creates synthetic image-question pairs for a minimal-context tasks to probe three distinct facets of nonlocal visual reasoning: comparative perception, saccadic search, and smooth visual search. Our generator, evaluation sets, and evaluation code are available here.
  • We conduct a comprehensive evaluation of leading VLMs (including Gemini 2.5 Pro, Claude Vision 3.7, GPT-o3, and others) demonstrating that even flagship models lag far behind humans on trivial visual reasoning tasks, despite advances in primitive vision.
  • We create several variants of each task to determine why models fail to perform as well as humans. We also determine under which circumstances they look carefully images and use their perception abilities.
Refer to captionRefer to caption
Figure 1: Object Re-Identification, top-left. The task refers to two images, marked ‘Image 1’ and ‘Image 2’ in the figure. The goal is to predict if the same object appears in the second image, allowing for a rigid transformation of the entire object. The first row shows the exact same image twice, with a 50% chance of a component shape shifting in the second image. The second row tests understanding of a rigid transformation of the entire object. Both of these require the ability to internalize the concept of an object that is not just a collection of pixels, but also has higher-level semantic meaning. Visual Scavenger Hunt, right. The goal is to follow the text written on the shape for a specified number of steps and state the color of the final shape. This requires the ability to “visually search" through the image and continue probing it for information based on what the model has already found in the image. For a human, this task is trivial. Circuit Connections, bottom-left. The goal is to follow the wire out of a specific port on the central breadboard.
@@ 2 Related Work
Benchmarks on Perception Primitives. VLMs achieve strong performance on many complex tasks, including OCR, image captioning, and scene understanding [21, 3, 6, 23, 1, 18]. However, this high-level competence contrasts with known deficiencies in low-level perception. For instance, VLMs can struggle with recognizing basic shapes or performing simple visual arithmetic [19, 9, 22]. Research suggests these perceptual limitations may originate in the language decoder, even with adequate image encoder representations [9, 19]. These foundational gaps raise questions about whether VLMs’ success always stems from robust visual processing, and critically, when and how they engage in visual algorithms. While benchmark suites like VisOnlyQA and HallusionBench evaluate specific visual weaknesses, such as hallucination and illusion failures under controlled settings [12, 7], we evaluate whether learned priors interfere not only with perception but also with visual reasoning.
Chart and Graph Understanding. The importance of visual data interpretation has motivated numerous evaluations and benchmarks such as ChartQA and MultiChartQA, which expose VLMs to diverse charts [15, 24, 11]. This area has also spurred the development of several models specifically trained for chart understanding [15, 14]. However, VLMs’ underwhelming performance on more recent benchmarks such as ChartQAPro suggests that they have yet to develop robust graph-understanding capabilities [16].
Visual Representations and Reasoning. Transformers are able learn representations that are robust to viewpoint, lighting, and occlusions [10, 17, 2, 8]. Nevertheless, VLMs struggle with visual representations outside of specific environments environments. Other diagnostic benchmarks reveal deficiencies in multi-view and multi-instance consistency, despite strong feature extraction capabilities [5]. These inconsistencies suggest that robust feature extraction alone does not guarantee human-like performance for image understanding. Furthermore, other research has shown that VLMs do not adequately model causal or physical relationships [20]. Such high-level reasoning failures may stem partly from models forming conclusions based on brief visual assessments, relying more on their language understanding than on thorough visual processing.
These evaluations pinpoint what VLMs fail at, but they do not provide a controlled setting to investigate specific reasoning patterns. Our synthetic evaluation isolates the parts of reasoning that occur in the visual domain and not in natural language.
@@ 3 Evaluation Design
Our evaluation investigates capabilities that are critical to achieving generalist vision.
The first capability is comparative perception, and is tested in our first task. Humans possess a remarkable ability to recognize objects with in different positions or with partial occlusion. This robustness allows us to identify the same object or distinguish between very similar ones across diverse settings[4]. This can be accomplished by a sequential comparison between two images to verify shared properties between them that are difficult to verbalize. For example, it is difficult to describe the differences between two faces. This ability hinges on what we call comparative perception, which involves a qualitative comparison between two images, one of which is in working memory.
The second ability is saccadic search, and is tested in our second task, Visual Scavenger Hunt. Updating a world model from an image requires using evidence gathered from multiple regions in the picture that have logical dependencies with each. Saccadic search is the ability to continually analyze an image conditioned on past information gleaned from that same image.
The third ability is smooth visual search, and is the ability humans have to track an object in motion. It is also used when following lines or paths. This task is purely visual and cannot be effectively simulated in natural language.
@@
We use the Object Re-Identification task to test comparative perception: the model must hold two views in working memory and compare them under allowed transformations. Each instance of the task involves two images, ‘Image 1’ and ‘Image 2’. 2–6 shapes are shown in ‘Image 1’. These shapes undergo a random entire-object-level rigid transformation in ‘Image 2’, and are sometimes jittered. The task is to distinguish the pairs of images that are identical after rigid transformation from those that are not. We render distractor objects in ‘Image 2’ to reflect the human vision capacity to re-identify objects across different environments. Examples of this task can be found in Figure 1.
Our goal is to test re-identification capabilities that are trivial for humans. To ensure our evaluation is robust and truly assesses this ability, we enforce specific generation parameters. We avoid nearly imperceptible transformations, such as small rotations, translations, and re-scalings, as we do not aim to test pixel-level accuracy. Furthermore, the object is always kept in-frame and unoccluded.
To disentangle where a VLM might succeed or fail, we present three presentation variants. In the Standard variant, all component parts are physically contiguous. It is designed to test the most intuitive concept of what an “object" is. In the Unconnected variant, we generate the objects randomly and do not require them to be connected. This is designed to evaluate a more abstract concept of an object, and also determine how connectedness affects VLM perception. In the Pixel-Perfect variant, we do not apply any rigid transformation, meaning that in positive examples ‘Image 2’ is an exact pixel-for-pixel match of ‘Image 1’ (except for the distractor shapes). In negative examples, one or more of the component shapes is still jittered. Thus, this trial is solvable by direct pixel-matching. Examples of all three tasks can be found in the Appendix.
The examples are generated from a uniform distribution between “Yes" and “No", so the random guessing baseline is 50%.
@@
This task tests saccadic search: saccadic search: the ability of the model to make discrete, evidence-driven jumps across the image. In many real-world visual challenges, locating a pixel cluster with semantic content is not enough. Instead, the task adapts as the agent gathers new clues. For example, no flame on a gas stove does not signal a gas leak—unless the knob is on. Humans use each observation to decide where to look next, and this capability is vital to generalist agents.
We design our Visual Scavenger Hunt task to evaluate iterative visual search explicitly. The model is presented with a grid of different-colored shapes, each labeled with a different color-shape pair. The prompt provides a starting shape, and requires the model to follow a straightforward “scavenger hunt" of shapes around the image for a specified number of steps (which we call the chain length). The scene is randomly generated, so the entire scene must be searched to find a specific shape. A visual example can be found in Figure 1. We confirm before testing that Qwen-2.5-7B and LLaMA-3.2-11B are both able to read the overlaid text and identify individual shapes at the rendered sizes, reducing the challenge to one of pure visual reasoning. The prompts that we used to do so can be found in the Appendix.
@@
We frame Circuit Connections as an instance of smooth visual search. The model is shown a procedurally generated circuit diagram with a breadboard, several components, and wires. The task is to specify which component a breadboard port is connected to; the model must trace a continuous contour—a wire—from its source to its terminal point. An example of this appears in Figure 1 (bottom).
Inspired by the “Subway Connections" task in BlindTest[19], we refine the task to directly evaluate the ability to follow the wire. Our synthetic generator contains three variants: the Standard, Single Color, and Unique Colors versions. On the Standard trial, each wire is one of five colors, selected randomly. The Unique Colors variant is designed as a control; each wire has a unique color so that the task is solvable by only looking at the ends of the wire. In the Single Color examples, all wires in an image have the same color. This trial is designed to prevent the model from taking shortcuts—by associating colors with locations or reasoning in natural language—rather than following the wire manually.
@@
The rendered image contains between 4-10 components, chosen from a uniform distribution. Thus, a random guessing baseline is approximately 14.29%.
@@ 4 Experiments
We evaluate Gemini 2.5 Pro, Claude-Vision 3.7, o4-mini and many open-source models on our benchmark in a few-shot setting. o3 is also included on the main trial for each task. All error bars represent standard error. Full evaluation details, including prompts, can be found in the Appendix.
Before running our experiments, we manually self-evaluated each task’s main category to find a human baseline. Our evaluator scored 100% on the Object Re-Identification and Visual Scavenger Hunt tasks and 99.5% on the Circuit Connections trial.
@@ 4.1 Object Re-Identification
The results for Object Re-Identification are summarized in Figure 2. On the Default variant, no model significantly outperforms random chance, indicating that they cannot perform comparative perception. However, on the other two trials, o4-mini, Gemini 2.5 Pro, and Claude 3.7 Sonnet improve by between 12-17%. They remain over 25 percentage points below the human baseline in all trials.
The F1+ and F1-scores, displayed in Figure 3, subdivide the models into three classes. Firstly, models that essentially ignore the input and almost always predict the same result (Molmo 7b and Llama 3.2 11b). The vast majority of models attempt to answer the question and are not especially biased across all trials, but give poor predictions for both classes. The third class consists of models that improve in the latter two variants (o4-mini, Gemini 2.5 Pro and Claude 3.7 Sonnet).
@@
The stronger VLMs display two separate modes of behavior. They perform similarly to all other models in the default setting but improve drastically on the other two variants.
The improved performance on the Unconnected variant compared to the Standard (connected) variant suggests that VLMs struggle with intra-component analysis but are more successful at identifying relationships between entities. Many models, especially smaller architectures, are constrained by core visual processing limitations; their failure to surpass random chance on any variant indicates an inability to perform basic visual comparisons. In contrast, higher-performing models possess higher visual capacity but only utilize it selectively. They do not perform well when objects are presented as coherent, connected entities. These models have strong implicit priors about entity analysis.
Figure 3: F1 for positive and negative classes across all trials of Object Re-Identification. Some models predict identically across the entire trial (red box.) The strong models (orange box) perform poorly on the standard variant, but become better at recognizing similar objects in the other two trials.
@@
These models also perform about equally well on the Pixel-Perfect evaluation. Gemini-2.5-Pro scored 100% accuracy on the 11 responses in which it provided more than 10 tokens. We theorize that because the first and second image are so similar, these problems are easier to convert to an instance of natural-language comparison as any perturbation disqualifies the second object from being identical to the first.
These results indicate that VLMs are currently incapable of general comparative perception and do not always examine the images carefully. Even when they look closely, their vision is much less robust than that of humans.

### v2

@@ Abstract
Vision-Language Models (VLMs) excel at complex visual tasks such as VQA and chart understanding, yet recent work suggests they struggle with simple perceptual tests. We present an evaluation that tests vision-language models’ capacity for nonlocal visual reasoning—reasoning that requires chaining evidence collected from multiple, possibly distant, regions of an image. We isolate three distinct forms of nonlocal vision: comparative perception, which demands holding two images in working memory and comparing them; saccadic search, which requires making discrete, evidence‑driven jumps to locate successive targets; and smooth visual search, which involves searching smoothly along a continuous contour. Flagship models (e.g. GPT-5, Gemini 2.5 Pro, Claude Sonnet 4), even those that perform well on prior primitive‑vision benchmarks, fail these tests and barely exceed random accuracy on two variants of our tasks that are trivial for humans. Our structured evaluation suite allows us to test if VLMs can perform similar visual algorithms to humans. Our findings show that despite gains in raw visual acuity, current models lack core visual reasoning capabilities.
@@ 1 Introduction
Vision-Language Models (VLMs) have demonstrated impressive performance on complex multimodal tasks. These models appear to possess a deep understanding of images and achieve over 90% accuracy on benchmarks like AI2D and ChartQA [kembhavi2016diagrams, Masry2023UniChart, masry2022chartqabenchmarkquestionanswering]. Such high-level competence would suggest these models have strong primitive visual skills. However, recent work such as VLMs are Blind [Rahmanzadehgervi2024Blind] or HallusionBench [Guan2024HallusionBench], shows that VLMs excel at answering high-dimensional questions that require background context but fail at recognizing simple geometry. Yet newer models score higher on these adversarial benchmarks, indicating that they possess stronger visual acuity.
@@
This strategy works only when combined with known assumptions about an image. Graph comprehension questions are embedded with strong priors such as standardized layouts (e.g., axes, legends) and may not require the model to reason over nonlocal regions. Models could instead learn to exploit these learned commonalities and only superficially parse graphs. This reliance on convention over direct visual evidence would stunt models from parsing images that defy these norms. Indeed, recent literature supports the hypothesis that VLMs are not as proficient at reading graphs as benchmark scores suggest [Masry2025ChartQAPro]. For example, Masry et al. [Masry2025ChartQAPro] showed that top models’ performance catastrophically declined on their novel ChartQAPro dataset compared to ChartQA, indicating a missing skill set for handling diverse, unseen charts. This, along with findings from benchmarks like HallusionBench [Guan2024HallusionBench], suggests VLMs still prioritize background knowledge over the visual evidence presented to them.
We identify three core types of nonlocal visual reasoning. First, comparative perception is the qualitative comparison of visual entities even when precise discrepancies are difficult to articulate (e.g., recognizing that two complex shapes are not identical without explicitly itemizing every differing feature). Second, saccadic search, named after the rapid eye movements humans make, is the process of gathering and integrating information from different image regions. Each piece of evidence informs the next step; for example, this process is used when consulting a chart’s legend, then locating the corresponding line, then referencing an axis. Third, smooth visual search describes the continuous tracing of visual elements, such as following the outline of an object or tracing a curve to its conclusion. To systematically evaluate these capabilities, we introduce a procedurally-generated evaluation set comprising three task categories designed to be trivial for humans and require minimal prior knowledge. We present three task categories: Object Re-Identification, Visual Scavenger Hunt, and Circuit Connections. Examples of these tasks appear in Figure 1.
@@
    When do VLMs err in basic perception or perceive correctly but fail at visual reasoning? Do early perceptual mistakes compound or self-correct during reasoning?
@@
    Can VLMs perform comparative perception and saccadic search? If so, must these models use natural language judgments to guide these processes, or can they execute these tasks through direct visual analysis?
@@
All tested models perform poorly on at least one variant of a task that is trivial for humans. The performance gradient on Object Re-Identification shows that modern models selectively choose when to examine an image closely. Our work further shows that fuzzy vision interferes with visual reasoning and the models seem unable to self-correct. They instead rely on their prior natural language judgments over direct evidence in the image. VLMs struggle the most with smooth visual search and our analysis of when they succeed suggests that most models cannot trace lines.
@@
  • We release three procedural generators for an evaluation suite. Each generator creates synthetic image-question pairs for a minimal-context task to probe three distinct facets of nonlocal visual reasoning: comparative perception, saccadic search, and smooth visual search. Our generator, evaluation sets, and evaluation code are available here.
  • We conduct a comprehensive evaluation of leading VLMs (including GPT-5, Gemini 2.5 Pro, and Claude Sonnet 4) demonstrating that even flagship models lag far behind humans on trivial visual reasoning tasks, despite advances in primitive vision.
  • We create several variants of each task to determine why models fail to perform as well as humans. We also determine under which circumstances they look carefully at images and use their perception abilities.
Figure 1: Object Re-Identification (top): Determine whether the same object that appears in Image 1 also appears in Image 2, up to a transform of the entire object but not individual component shapes. Visual Scavenger Hunt (bottom-left): From the indicated shape, follow the written labels for the specified count and report the final shape’s color. Circuit Connections (bottom-right): From a named port on the central breadboard, trace the wire to its endpoint. The prompts here are abbreviated; full instructions are in the appendix.
@@ 2 Related Work
Benchmarks on Perception Primitives. VLMs achieve strong performance on many complex tasks, including OCR, image captioning, and scene understanding [Touvron2023LLaMA, Bai2025Qwen25, Deitke2024Molmo, Zhu2025InternVL3, Abouelenin2025Phi4Mini, Mistral2025Small3.1]. However, this high-level competence contrasts with known deficiencies in low-level perception. For instance, VLMs can struggle with recognizing basic shapes or performing simple visual arithmetic [Rahmanzadehgervi2024Blind, Huang2025CogAlign, Wei2024SlowPerception]. Research suggests these perceptual limitations may originate in the language decoder, even with adequate image encoder representations [Huang2025CogAlign, Rahmanzadehgervi2024Blind]. These foundational gaps raise questions about whether VLMs’ success always stems from robust visual processing, and critically, when and how they engage in visual algorithms.
Benchmarks on Visual Reasoning. To probe the underlying algorithmic visual skills necessary for robust image understanding, we take a different approach from existing visual reasoning benchmarks. Evaluations like Bongard problems and ARC [bongard1970pattern, chollet2019measureintelligence] test complex reasoning and are constructed as abstract puzzles that are challenging humans. Failing such a task does not distinguish between a failure in high-level abstract reasoning and a failure in fundamental visual processing. Our evaluation isolates these primitive visual skills by stripping away the puzzle-solving layer. While benchmark suites like VisOnlyQA and HallusionBench evaluate specific visual weaknesses, such as hallucination and illusion failures under controlled or adversarial settings [Kamoi2024VisOnlyQA, Guan2024HallusionBench], we evaluate general visual algorithms.
Chart and Graph Understanding. The importance of visual data interpretation has motivated numerous evaluations and benchmarks such as ChartQA and MultiChartQA, which expose VLMs to diverse charts [masry2022chartqabenchmarkquestionanswering, Zhu2024MultiChartQA, Islam2024ChartComprehension]. This area has also spurred the development of several models specifically trained for chart understanding [Masry2023UniChart, lee2023pix2struct]. However, VLMs’ underwhelming performance on more recent benchmarks such as ChartQAPro suggests that they have yet to develop robust graph-understanding capabilities [Masry2025ChartQAPro].
@@ 3 Evaluation Design
Our evaluation isolates three core capabilities of nonlocal visual reasoning, each grounded in principles from classical human vision research. The first, comparative perception, draws on work in perceptual organization and attention. Classic Gestalt principles such as proximity and connectedness suggest that humans automatically group visual elements into coherent objects [wertheimer1938gestalt], while Treisman’s Feature Integration Theory (FIT) [treisman1980feature] shows that attention serves to bind distributed features into a unified representation. By comparing Object Re-Identification performance on the Unconnected variant against the Standard variant (where components are contiguous), we directly test whether models are sensitive to the Gestalt principle of connectedness. The second capability, saccadic search, is motivated by Ullman’s theory of visual routines [ullman1984visual], which proposed that complex visual operations are composed of primitive procedures such as shifting the focus of attention and indexing marked locations. Specifically, it is based on systematic scanning, or using each observation to guide the next. The third capability, smooth visual search, corresponds to the visual routine of boundary and contour tracing within Ullman’s framework.
@@
We use this task to test comparative perception: the model must hold two views in working memory and compare them under allowed transformations. Each instance of the task involves two images, ‘Image 1’ and ‘Image 2’. 2–6 shapes are shown in ‘Image 1’. In ‘Image 2’, this entire object undergoes a random transformation (i.e., rotation, translation, and scaling) and is always kept fully in-frame.
In half of the examples (the negative cases), one or more of the component shapes are also transformed independently to create a structurally different object. We avoid nearly imperceptible transformations, such as small rotations, translations, and re-scalings, as we do not aim to test pixel-level accuracy. Additionally, we render distractor objects in ‘Image 2’ to reflect the human capacity to re-identify objects across different environments. We restrict these distractor objects from occluding any components of the original object. Examples of this task can be found in Figure 1.
To disentangle VLM success conditions, we present three presentation variants. The Standard variant renders objects with physically contiguous component shapes to test the most intuitive concept of what an object is. The Unconnected variant removes this contiguity requirement to evaluate a more abstract object concept and determine how connectedness affects VLM perception. In the Pixel-Perfect variant, we do not apply any transformation to the object as a whole. Thus, positive examples of ‘Image 2’ are pixel-for-pixel matches of ‘Image 1’ (except for the added distractors). Examples of all three tasks can be found in the Appendix.
The examples are generated from a uniform distribution between “Yes” and “No”, so the random guessing baseline is 50%.
@@
This task tests saccadic search: the ability of the model to make discrete, evidence-driven jumps across the image. In many real-world visual challenges, locating a pixel cluster with semantic content is not enough. Instead, the task adapts as the agent gathers new clues. For example, an unlit gas stove does not signal a gas leak unless the knob is in the on position. Humans use each observation to decide where to look next, and this capability is vital to generalist agents.
We design our Visual Scavenger Hunt task to evaluate iterative visual search explicitly. The model is presented with a grid of different-colored shapes, each labeled with a different color-shape pair. The prompt provides a starting shape, and requires the model to follow a straightforward “scavenger hunt” of shapes around the image for a specified number of steps (which we call the chain length). The scene is randomly generated, so the entire scene must be searched to find a specific shape. A visual example can be found in Figure 1.
@@
We frame Circuit Connections as an instance of smooth visual search. The model is shown a procedurally generated circuit diagram with a breadboard, several components, and wires. The task is to specify which component a breadboard port is connected to; the model must trace a continuous contour—a wire—from its source to its terminal point. An example of this appears in Figure 1.
Inspired by the “Subway Connection” task in BlindTest [Rahmanzadehgervi2024Blind], we refine the task to directly evaluate the ability to follow the wire. Our synthetic generator contains three variants: the Standard, Single Color, and Unique Colors versions. On the Standard trial, each wire is one of five colors, selected randomly. The Unique Colors variant is designed as a control; each wire has a unique color so that the task is solvable by only looking at the ends of the wire. In the Single Color examples, all wires in an image have the same color. This trial is designed to prevent the model from taking shortcuts—by associating colors with locations or reasoning in natural language—rather than following the wire manually.
@@
The rendered image contains between 4-10 components, chosen from a uniform distribution. Thus, a random guessing baseline is 14.29%.
@@ 4 Experiments
We evaluate Gemini 2.5 Pro, Claude Sonnet 4, GPT-5, as well as other closed and open-source models on our benchmark in a few-shot setting. All error bars represent standard error. The full evaluation details, including prompts, can be found in the Appendix.
Before running our experiments, we manually self-evaluated 200 examples of each task’s main category to find a human baseline. Our evaluators scored 100% on the Object Re-Identification and Visual Scavenger Hunt tasks and 99.5% on the Circuit Connections trial.
@@ 4.1 Object Re-Identification
The results for Object Re-Identification are summarized in Figure 2. On the Standard variant, no model significantly outperforms random chance, indicating that they cannot perform comparative perception. However, on the other two trials, models such as GPT-5, Gemini 2.5 Pro, and Claude Sonnet 4 score between 12-24 percentage points higher. They remain over 20 percentage points below the human baseline in all trials.
The F1 (positive) and F1 (negative) scores, displayed in Figure 3, subdivide the models into three classes. Firstly, models that essentially ignore the input and almost always predict the same result (Molmo 7b and Llama 3.2 11b). The vast majority of models attempt to answer the question and are not especially biased across all trials, but give poor predictions for both classes. The third class consists of models that improve in the latter two variants.
@@
A group of models—Molmo (7B), Llama-3.2 Vision (11B), and Phi-4 Multimodal (14B)— effectively ignore the input and almost always predict the same result, as evidenced by their F1 scores being 0.00 for one class across multiple variants (Figure 3). This indicates they either cannot perform comparative perception or do not attempt to do so across all trials.
The models highlighted in the orange box in Figure 3 fail in ways that show they have the ability to selectively compare objects only in certain circumstances. On the Standard variant, their accuracy is near random chance, with the highest score being 60%. However, their performance significantly improves on both the Unconnected (GPT-5: 77%, Gemini 2.5 Pro: 65%, Claude Sonnet 4: 79%) and Pixel-Perfect ( GPT-5: 71%, Gemini 2.5 Pro: 65%, Claude Sonnet 4: 70%) variants. In the Standard variant, the component shapes of the object are always physically contiguous. According to Treisman’s Feature Integration Theory, humans bind basic features—color, shape, orientation—into unified object representations through focused attention. These models’ lower accuracy on the Standard variant indicates they do not bind objects the same way humans do. More specifically, their poor performance (<60%) suggests that these models struggle with fine-grained inspection of connected entities. This contrasts with human perception, where Gestalt principles of grouping (e.g., by proximity and connectedness) make connected objects easier to process and compare. Because the Standard variant is a subset of the Unconnected variant, we suspect that these models are capable of comparing different objects but engage this ability selectively. If two Gestalt-grouped objects are similar enough the model does not inspect them closely. They appear to bypass comparative attention when objects appear similar enough under Gestalt grouping cues.
Figure 3: F1 for positive and negative classes across all trials of Object Re-Identification. Some models predict identically across the majority of trials (red box.) The strong models (orange box) perform poorly on the standard variant, but become better at recognizing similar objects when tested on the other two trials.
@@
The stronger models also perform slightly worse on the Pixel-Perfect as they do on Unconnected. On this trial, Gemini-2.5-Pro scored 100% accuracy on the 11 responses in which it provided more than 10 tokens. We theorize that because the first and second images are so similar, these problems are easier to convert to an instance of natural-language comparison as any perturbation disqualifies the second object from being identical to the first.
Taken together, these results indicate that VLMs are currently incapable of generalist comparative perception and do not always examine the images carefully.